# FedCC: Robust Federated Learning against Model Poisoning Attacks

### v3

@@ Abstract
Federated learning is a distributed framework designed to address privacy concerns. However, it introduces new attack surfaces, which are especially prone when data is non-Independently and Identically Distributed. Existing approaches fail to effectively mitigate the malicious influence in this setting; previous approaches often tackle non-IID data and poisoning attacks separately. To address both challenges simultaneously, we present FedCC, a simple yet effective novel defense algorithm against model poisoning attacks. It leverages the Centered Kernel Alignment similarity of Penultimate Layer Representations for clustering, allowing the identification and filtration of malicious clients, even in non-IID data settings. The penultimate layer representations are meaningful since the later layers are more sensitive to local data distributions, which allows better detection of malicious clients. The sophisticated utilization of layer-wise Centered Kernel Alignment similarity allows attack mitigation while leveraging useful knowledge obtained. Our extensive experiments demonstrate the effectiveness of FedCC in mitigating both untargeted model poisoning and targeted backdoor attacks. Compared to existing outlier detection-based and first-order statistics-based methods, FedCC consistently reduces attack confidence to zero. Specifically, it significantly minimizes the average degradation of global performance by 65.5%. We believe that this new perspective on aggregation makes it a valuable contribution to the field of FL model security and privacy. The code will be made available upon acceptance.
@@ 1 Introduction
Federated Learning (FL) [22] is a distributed model training framework designed to preserve privacy. FL allows data to remain on client devices but exchanges their local model parameters only for model training. It thus reduces (data) privacy risks associated with centralized learning, not requiring central data collection. Thus, FL is especially suitable for data-sensitive environments since raw data remains on clients, reducing the risk of data leakage on the server or in between communication.
@@
Some preliminary defenses propose robust aggregation methods, but they may compromise privacy by sharing raw data or exposing data distribution to untrusted parties [6, 30]. While some defenses exist, most focus on IID data settings, which are rare in real-world scenarios. Non-IID data, with variations in class distribution and features across clients [12], can gradually degrade model performance and complicate the detection of malicious clients [39, 35], as the impact of attacks increases with the degree of non-IID [19, 26].
In this study, we introduce FedCC, a defense mechanism that mitigates untargeted model poisoning and targeted backdoor attacks in both IID and non-IID settings. FedCC leverages the Centered Kernel Alignment (CKA) of Penultimate Layer Representations (PLRs) to distinguish malicious clients from benign ones. PLRs are known to be highly specific to local data distributions [36], making clients distinguishable, especially when data is non-IID. Exploiting CKA’s well-known capability of measuring similarity in high-dimensional space, such as neural network representations, we accurately measure the similarity between the PLR of different clients, further allowing more accurate detection of malicious clients. Empirical evidence supporting our rationale is shown in Figure 2 and Table 1. The use of PLRs shows the highest separability between benign and malicious clients in all settings, as seen by the largest distance between clusters in the PLR. The use of CKA, specifically kernel CKA, is justified by its highest performance in both targeted and untargeted IID and non-IID settings. Furthermore, our method maintains data privacy by using similarity measures between local and global models, not having shared data. We named our proposed method FedCC, Federated learning using Clustered CKA.
Empirical experiments show that FedCC outperforms existing defenses that rely on first-order statistics, which usually suffer from a high false negative detection rate. It significantly improves global accuracy against untargeted attacks (66.27% compared to similar approaches) and reduces attack confidence in targeted backdoor attacks to almost zero while preserving main task accuracy. These results highlight FedCC’s effectiveness and robustness in mitigating attacks in FL settings, particularly in non-IID data settings.
@@
  • We propose FedCC, a novel and scalable defense method against model poisoning attacks using the CKA of PLRs of local and global models and performing layer-wise weighted aggregation of their model parameters.
  • We justify that PLR is a highly distinguishable feature to detect malicious clients by measuring the discrepancy between two clusters of clients.
  • We demonstrate that CKA provides the most sensitive and accurate similarity measure between differently trained models when the server has no knowledge about clients or centralized data, even in non-IID settings.
  • We empirically demonstrate the effectiveness of FedCC through exhaustive experiments. Our results show that comparing the CKA of PLRs filters malicious clients more reliably than the first-order statistics. FedCC outperforms existing defenses even in non-IID settings.
@@ 2.3 Penultimate Layer Representation in FL
The penultimate layer is a neural network model’s second-to-last layer (i.e., before the softmax layer). Wang et al. [30] discovered that PLRs are effective in distinguishing malicious models from benign ones; within benign clients, benign PLRs follow the same distribution, whereas malicious PLRs do not, across datasets and neural network architectures. Specifically, they demonstrated that distances among benign PLRs are smaller than those of benign and malicious PLRs. Their proposed method, FLARE, assigned a trust score to each client based on pairwise PLR discrepancies defined by Maximum Mean Discrepancy among all model updates, allocating lower values to those farther from the benign distribution. The model updates are then scaled and averaged, weighted by each client’s trust score. Meanwhile, FLARE continuously redirects the global model using server-owned raw data, increasing the risk of data leakage and jeopardizing data privacy.
@@ 2.4 Centered Kernel Alignment
Centered Kernel Alignment (CKA) [13] is a highly accurate similarity metric that measures how similar two differently initialized or trained neural networks are, producing a value between 0 (not similar) and 1 (identical). CKA can be calculated by: $CKA(X,Y)=\frac{HSIC(X,Y)}{\sqrt{HSIC(X,X)HSIC(Y,Y)}}=\frac{{\lVert K(X^{T})K(Y)\rVert}^{2}}{{\lVert K(X^{T})K(X)\rVert}{\lVert K(Y^{T})K(Y)\rVert}}$ where $HSIC$ is Hilbert-Schmidt Independence Criterion [9] (i.e., a not normalized version of CKA) and $K$ is the RBF kernel. It constructs similarity kernel matrices in the weight space and compares these matrices to characterize the representation space, allowing comparisons between layers with varying widths and different initialization schemes.
CKA includes several vital properties for similarity measures of neural networks. It is non-invariant to invertible linear transformations, meaning similarity scores change if such transformations are applied. This property is vital; the gradient descent algorithm is not invariant to these transformations because similarity metrics invariant to linear transformations are inaccurate on models trained with gradient descent. It is also invariant to orthogonal transformations and isotropic scaling, which are desirable for neural networks trained with gradient descent owing to their stochastic nature. These properties make CKA a precise and reliable metric for measuring similarities between neural networks, particularly when comparing potential models from malicious clients.
CKA has been utilized to address data heterogeneity issues, taking into account that the similarity of non-IID data is notably lower than that of IID data [28], particularly in specific layers [27]. However, CKA has yet to be leveraged as a defense mechanism against attacks involving models from malicious clients where the data are distributed across the clients in a non-IID manner.
@@ 3.2 Attackers’ Capability
  • An attacker has knowledge about compromised clients’ data, such as current local model, previous global model, and hyperparameters.
@@ 3.3.1 Untargeted Model Poisoning Attacks [7]
Untargeted-Med [7] manipulates the model parameters based on the maximum and minimum so that chosen coordinate-wise median values direct toward an inverse direction. The attack starts with defining $w_{max,j}=\text{max}\{w_{(m+1),j},w_{(m+2),j},...,w_{n,j}\}$, which indicates the maximum, and $w_{min,j}=\text{min}\{w_{(m+1),j},w_{(m+2),j},...,w_{n,j}\}$ which indicates the minimum of the $j$th local model parameters on the benign clients, respectively. Also, to avoid sampled $m$ numbers being outliers, if $s_{j}=-1$, $m$ numbers in [$w_{max,j},b\cdot w_{max,j}$] (when $w_{max,j}>0$) or [$w_{max,j},b/w_{max,j}$] (when $w_{max,j}\leq 0$) are randomly sampled. Otherwise, $m$ numbers [$w_{min,j}/b,w_{min,j}$] (when $w_{min,j}>0$) or [$b\cdot w_{min,j},w_{min,j}$] (when $w_{min,j}\leq 0$) are randomly sampled. We set $b=2$ as the same as the paper [7].
@@ 3.3.2 Targeted Backdoor Attacks
are based on [3]. Each malicious client owns one sample of mislabeled images and trains the local model on it. We trained the model on backdoor tasks along with the main task(s); the training went on for both malicious and benign tasks to maintain benign accuracy such that backdoor training remains stealthy. Then we boosted malicious clients’ updates to negate the combined effect of the benign agent: $w_{i}^{t}=w_{G}^{t-1}+\alpha_{m}(w_{i}^{t-1}-w_{G}^{t-1})\text{ for }i=1,...,m$ where $t$ is the current epoch and $\alpha_{m}$ is a boosting factor. For DBA [33], we implant split backdoor triggers into the input so that the global model can be attacked in a distributed manner. For more detail, refer to [33]. In CIFAR10, for example, we manipulate the model parameters to misclassify an image of ‘airplanes’ as ‘bird’ whereas normally classifying the image of other classes.
@@ 4.1 Overview of FedCC
Figure 1 depicts an overview of FedCC. (1) A server initializes and advertises the global model to $n$ participating clients, selected out of total $k$ clients with a probability of fraction $C$. (2) The clients individually train their local models using their respective datasets and send the trained weights to the server for aggregation. Our proposed method intervenes at this moment. (3) The server extracts PLRs from both the global model and each local model, and (4) compares RBF Kernel CKA values between them. (5) The CKA values are then clustered into two, potentially representing benign and malicious clients. (6) The CKA similarity values are then normalized across all clients for every layer as $a$, but the second last layer’s CKA similarity values are normalized within the larger cluster as $b$, backed by the assumption that the number of malicious clients does not exceed $n/2-1$. (7) Finally, the server aggregates each layer of local models weighted by $a$, but the PLR weighted by $b$, to update a global model for the current epoch and proceeds to the next epoch by distributing the updated global model to the newly selected $n$ clients.
@@ 4.2 Detailed Design
A major challenge when filtering malicious clients in the presence of non-IID data is the server’s lack of knowledge regarding the data distribution. Even if the data is benign, local learning models can exhibit significant angular or magnitude differences if they are non-IID. Thus, we require a similarity metric independent of data distribution and not influenced by the distance or direction of model parameters.
@@ 4.2.1 Penultimate Layer
is the output of the second-to-last layer before the softmax layer in CNN. In federated learning, where the server cannot access local training data, it receives trained model parameters (weights) from selected clients. Therefore, the collected weights are the only available information for determining whether a client is malicious. PLR takes the end result of the convolution and pooling process that will reach a classification decision. Thus, we assumed that computed weights of the second last layer might contain the essential information that strongly influences classification results.
@@
To justify the utilization of PLRs, we measure the distance between two clusters of clients (benign and malicious) using dendrograms with a single linkage metric and correlation distance method. Figure 2 presents cluster distances of each layer under different attack scenarios, such as untargeted-mKrum attacks in IID or non-IID settings. Observably, the cluster distance is the most considerable in PLR to other layers, indicating the highest degree of differences. Based on this observation, we conclude that PLR contains the most distinguishable and indicative information. Therefore, we utilize the weights of the penultimate layer to compare their similarity to the global model.
4.2.2 Kernel CKA[13]
is a similarity metric to assess the similarity or dissimilarity between two neural networks. It compares representations at different layers or models trained in different ways. The sophisticated similarity comparison ability of Kernel CKA makes it a suitable choice for our purpose.
To justify the effectiveness of Kernel CKA, we compared its performance with other similarity metrics such as linear CKA, MMD (not normalized CKA), cosine similarity (angle), and Euclidean similarity (distance). We evaluated the test accuracy of FedCC using the CIFAR10 dataset under two untargeted attacks and one targeted attack, separately in IID and non-IID settings. The results are summarized in Table 1. In the ’targeted’ column, the values represent the combination of test accuracy and backdoor confidence, with higher test accuracy and lower backdoor confidence indicating better performance. We observed that Kernel CKA consistently yields the highest performance across experiments; therefore, we decided to utilize Kernel CKA to compare the similarities of PLRs.

### v4

@@ Abstract
Federated learning is a distributed framework designed to address privacy concerns. However, it introduces new attack surfaces, which are especially prone when data is non-Independently and Identically Distributed. Previous approaches often tackle non-IID data and poisoning attacks separately. To address both challenges simultaneously, we present FedCC, a simple yet effective novel defense algorithm against model poisoning attacks. It leverages the Centered Kernel Alignment similarity of Penultimate Layer Representations for clustering, allowing the identification and filtration of malicious clients, even in non-IID data settings. The penultimate layer representations are meaningful since the later layers are more sensitive to local data distributions, which allows better detection of malicious clients. The sophisticated utilization of layer-wise Centered Kernel Alignment similarity allows attack mitigation while leveraging useful knowledge obtained. Our extensive experiments demonstrate the effectiveness of FedCC in mitigating both untargeted model poisoning and targeted backdoor attacks. Compared to existing outlier detection-based and first-order statistics-based methods, FedCC consistently reduces attack confidence to zero. Specifically, it significantly minimizes the average degradation of global performance by 65.5%. We believe that this new perspective on aggregation makes it a valuable contribution to the field of FL model security and privacy. Code is available at https://github.com/HyejunJeong/FedCC.
@@ 1 Introduction
Federated Learning (FL) [23] is a distributed model training framework designed to preserve privacy by restricting data to remain on client devices. Only model parameters are exchanged, minimizing data privacy risks typically associated with centralized learning. This makes FL particularly useful in data-sensitive environments, where raw data should never leave clients, reducing the risk of data leakage.
@@
While some defenses rely on robust aggregation methods, many fail to maintain privacy by sharing raw data or exposing data distributions to untrusted parties [7, 31]. Furthermore, most existing defenses assume IID (Independent and Identically Distributed) data, which is rare in real-world scenarios. Non-IID data, where clients’ data distributions and features vary, complicates attack detection and performance maintenance [13, 39, 36]. As the degree of non-IID data increases, the impact of attacks also grows [20, 27].
In this study, we introduce FedCC, a defense mechanism against untargeted model poisoning and targeted backdoor attacks in both IID and non-IID settings. FedCC leverages Centered Kernel Alignment (CKA) of Penultimate Layer Representations (PLRs) to distinguish malicious clients from benign ones. PLRs are highly sensitive to local data distributions, making clients distinguishable, especially in non-IID environments. By exploiting CKA’s ability to measure similarity in high-dimensional spaces, we can accurately identify malicious clients. Empirical evidence, shown in Figure 3 and Table 1, demonstrates that PLRs provide the highest separability between benign and malicious clients, with CKA delivering superior performance in both IID and non-IID settings.
Importantly, FedCC ensures data privacy by relying on similarity measures between models instead of shared data. The proposed method outperforms existing defenses based on first-order statistics, which typically suffer from high false negative rates. Our experiments show that FedCC significantly improves global accuracy against untargeted attacks (66.27% improvement) and reduces attack confidence in targeted backdoor attacks to nearly zero while preserving main task accuracy.
@@
  • We propose FedCC, a novel and scalable defense method that uses CKA of PLRs and performs layer-wise weighted aggregation of model parameters to defend against model poisoning attacks.
  • We demonstrate that PLRs provide a highly distinguishable feature for detecting malicious clients by measuring the discrepancy between clusters of clients.
  • We justify the use of CKA as an accurate and sensitive similarity measure for comparing models, even in non-IID settings, when the server has no access to client data.
  • We empirically validate the effectiveness of FedCC through extensive experiments, showing that it outperforms existing defenses, especially in non-IID scenarios.
@@ 2.3 Penultimate Layer Representation in FL
Figure 1: LeNet (CNN) architecture and Penultimate Layer
The penultimate layer is a neural network model’s second-to-last layer (i.e., before the softmax layer), as shown in Figure 1. Wang et al. [31] discovered that PLRs are effective in distinguishing malicious models from benign ones; within benign clients, PLRs follow the same distribution, whereas malicious PLRs do not—across datasets and neural network architectures. Specifically, they demonstrated that distances among benign PLRs are smaller than those between benign and malicious PLRs. Their proposed method, FLARE, assigned a trust score to each client based on pairwise PLR discrepancies defined by Maximum Mean Discrepancy among all model updates, allocating lower values to those farther from the benign distribution. The model updates are then scaled and averaged, weighted by each client’s trust score. Meanwhile, FLARE continuously redirected the global model using server-owned raw data, increasing the risk of data leakage and jeopardizing data privacy.
@@ 2.4 Centered Kernel Alignment
CKA [14] is a highly accurate similarity metric used to measure how similar two differently initialized or trained neural networks are, producing a value between 0 (no similarity) and 1 (identical). It is computed as:
$CKA(X,Y)=\frac{HSIC(X,Y)}{\sqrt{HSIC(X,X)HSIC(Y,Y)}}=\frac{{\lVert K(X^{T})K(Y)\rVert}^{2}}{{\lVert K(X^{T})K(X)\rVert}{\lVert K(Y^{T})K(Y)\rVert}}$
where $HSIC$ refers to the Hilbert-Schmidt Independence Criterion [10], a non-normalized variant of CKA, and $K$ denotes the RBF kernel. CKA constructs similarity kernel matrices in the weight space and compares them to characterize the representation space, enabling comparison across layers with different widths and initialization schemes.
CKA satisfies several key properties desirable in neural network similarity metrics. It is non-invariant to invertible linear transformations, meaning similarity scores change if such transformations are applied. This property is vital; the gradient descent algorithm is not invariant to these transformations because similarity metrics invariant to linear transformations are inaccurate on models trained with gradient descent. It is also invariant to orthogonal transformations and isotropic scaling, both of which are desirable for neural networks trained with gradient descent owing to their stochastic nature. These properties make CKA a precise and reliable metric for measuring similarities between neural networks, particularly when comparing models potentially generated by malicious clients.
CKA has been utilized to address data heterogeneity issues, taking into account that the similarity of non-IID data is notably lower than that of IID data [29], especially in specific network layers [28]. However, it has not yet been explored as a defense mechanism against malicious clients, where model similarity could reveal suspicious deviations under non-IID distributions.
@@ 3.2 Attackers’ Capability
  • An attacker has knowledge about compromised clients’ data, such as the current local model, previous global model, and hyperparameters.
@@ 3.3.1 Untargeted Model Poisoning Attacks [8]
Untargeted-Med [8] manipulates the model parameters based on the maximum and minimum so that chosen coordinate-wise median values direct toward an inverse direction. The attack starts with defining the maximum and minimum of the $j$th local model parameters on the benign clients, $w_{max,j}=\text{max}\{w_{(m+1),j},w_{(m+2),j},...,w_{n,j}\}$ and $w_{min,j}=\text{min}\{w_{(m+1),j},w_{(m+2),j},...,w_{n,j}\}$, respectively. Also, to avoid sampled $m$ numbers being outliers, if $s_{j}=-1$, $m$ numbers in [$w_{max,j},b\cdot w_{max,j}$] (when $w_{max,j}>0$) or [$w_{max,j},b/w_{max,j}$] (when $w_{max,j}\leq 0$) are randomly sampled. Otherwise, $m$ numbers [$w_{min,j}/b,w_{min,j}$] (when $w_{min,j}>0$) or [$b\cdot w_{min,j},w_{min,j}$] (when $w_{min,j}\leq 0$) are randomly sampled. We set $b=2$ as the same as the paper [8].
@@ 3.3.2 Targeted Backdoor Attacks
are based on [3]. Each malicious client owns one sample of mislabeled images and trains the local model on it. We trained the model on backdoor tasks along with the main task(s); the training went on for both malicious and benign tasks to maintain benign accuracy such that backdoor training remains stealthy. Then we boosted malicious clients’ updates to negate the combined effect of the benign agent: $w_{i}^{t}=w_{G}^{t-1}+\alpha_{m}(w_{i}^{t-1}-w_{G}^{t-1})\text{ for }i=1,...,m$ where $t$ is the current epoch and $\alpha_{m}$ is a boosting factor. For DBA [34], we implant split backdoor triggers into the input so that the global model can be attacked in a distributed manner. For more details, refer to [34]. In CIFAR10, for example, we manipulate the model parameters to misclassify an image of ‘airplanes’ as ‘birds’, whereas normally it would classify the image as other classes.
@@ 4.1 Overview of FedCC
Figure 2 depicts an overview of FedCC. (1) The server initializes and broadcasts the global model to $n$ clients, selected from a total of $k$ candidates with selection fraction $C$. (2) Each client trains a local model on its private dataset and sends the updated weights to the server—this is where our proposed method intervenes. (3) The server extracts PLRs from both the global model and each local model, and (4) compares RBF Kernel CKA values between them. (5) The CKA values are then clustered into two, potentially representing benign and malicious clients. (6) CKA similarity values are normalized across all clients for each layer as $a$. For the second-to-last layer (PLR), normalization is performed within the larger cluster as $b$, under the assumption that malicious clients are fewer than $n/2-1$. (7) Finally, the server aggregates each layer of local models weighted by $a$, but the PLR weighted by $b$, to update a global model for the current epoch and proceeds to the next epoch by distributing the updated global model to the newly selected $n$ clients.
@@ 4.2 Detailed Design
A key challenge in filtering malicious clients under non-IID conditions is the server’s ignorance of the underlying data distribution. Even if the data is benign, local learning models can exhibit significant angular or magnitude differences if they are non-IID. Thus, we require a similarity metric independent of data distribution and not influenced by the distance or direction of model parameters.
@@ 4.2.1 Penultimate Layer
is the output of the second-to-last layer before the softmax layer in CNN. In federated learning, where the server cannot access local training data, it receives trained model parameters (weights) from selected clients. Thus, the received weights serve as the sole basis for detecting malicious behavior. The PLR captures the final representation produced by convolution and pooling before reaching a classification decision. We hypothesize that the PLR contains the most task-relevant, discriminative features.
@@
To test this, we measure the distance between two clusters of clients (benign and malicious) using dendrograms with a single linkage metric and correlation distance method. Figure 3 presents cluster distances of each layer under different attack scenarios, such as untargeted-mKrum attacks in IID or non-IID settings. Observably, the cluster distance is the most considerable in PLR to other layers, indicating the highest degree of differences. Based on this observation, we conclude that PLR contains the most distinguishable and indicative information. Therefore, we utilize the weights of the penultimate layer to compare their similarity to the global model.
@@
4.2.2 Kernel CKA[14]
is a similarity metric to assess the similarity or dissimilarity between two neural networks. It enables the comparison of representations across layers or among models trained under varying conditions. The sophisticated similarity comparison ability of Kernel CKA makes it suitable for our purpose.
To justify the effectiveness of Kernel CKA, we compared its performance with other similarity metrics such as linear CKA, MMD (not normalized CKA), cosine similarity (angle), and Euclidean similarity (distance). We evaluated the test accuracy of FedCC using the CIFAR10 dataset under two untargeted attacks and one targeted attack, separately in IID and non-IID settings. The results are summarized in Table 1. In the ‘targeted’ column, the values represent the combination of test accuracy and backdoor confidence, with higher test accuracy and lower backdoor confidence indicating better performance. We observed that Kernel CKA consistently yields the highest performance across experiments; therefore, we adopted it to measure PLR similarity.